# Efficient generation of divergent and collimated hot electrons via a novel multi-beam two-plasmon decay and stimulated Raman scattering mechanism

### v1

@@ Abstract
In inertial confinement fusion (ICF) implosions, the preheating risks associated with hot electrons generated by laser plasma instabilities (LPI) are contingent upon the angular characteristics of these hot electrons for a given total energy. Using particle-in-cell simulations, we reveal a novel multi-beam collaborative mechanism of two-plasmon decay (TPD) and stimulated Raman scattering (SRS), and investigate the angular variations of hot electrons generated from this TPD-SRS instability (TSI) driven collectively by dual laser beams with varying incident angles $\theta_{in}$ ($24^{\circ}$ to $55^{\circ}$ at the incident plane) for typical ICF conditions. In the simulations with $\theta_{in}\gtrsim 44^{\circ}$, TSI emerges as the dominant mechanism responsible for hot electron generation, leading to a wide angular distribution of hot electrons that exhibit both pronounced divergent and collimated components. The common Langmuir wave associated with TSI plays a crucial role in accelerating both components. By properly modeling the TSI common wave gains, we establish scaling relations between these gains and the energies of collimated and divergent hot electrons. These relations reveal that the divergent hot electrons are more sensitive to variations in gain compared to the collimated electrons. Additionally, the calculated gains qualitatively predict the asymmetry in hot electron angular distributions when the density gradients deviate from the bisector of the laser beams. Our findings offers insights for hot electron generation with multiple beams, potentially complementing previous experiments that underscore the critical role of overlapped intensity from symmetric beams within the same cone and the dominance of dual-beam coupling.
@@ I Introduction
In inertial confinement fusion (ICF), a spherical capsule is imploded by multiple high-power laser beams to compress the central fuels to extreme density and pressure, thereby initiating self-sustained fusion reactionsAtzeni and Meyer-ter Vehn (2004). During the implosion, hot electrons (with energy above $\sim 50$ keV) produced by laser plasma instabilities(LPI)W.L.Kruer (2003) can preheat the uncompressed fuel and degrade the implosionChristopherson et al. (2021); Solodov et al. (2022). In most ICF schemesNuckolls et al. (1972); He et al. (2016); Zhang et al. (2020a); Lan (2022), such hot electrons need to be reduced to improve the implosion, except for shock ignitionBetti et al. (2007) (SI), in which hot electrons generated during the spike pulse can enhance the shock pressureBaton et al. (2020); Tentori et al. (2021) and enhance the compression. While ignition has been achieved using the indirect-drive scheme on the NIF laser facilityAbu-shawareb (2024) with hot electron threats minimized by unique experimental configurations, the pursuit of higher energy gainsBatani et al. (2023); Sui and Lan (2024) and alternative ignition schemesGopalaswamy et al. (2024); Williams et al. (2024); He et al. (2016); Zhang et al. (2020a); Lan (2022); Betti et al. (2007) still demands effective control and thorough characterization of hot electron effects. Thus, reliably characterizing the properties of hot electrons is crucial to evaluate their potential detrimental or beneficial impacts.
Previous ICF experiments have explored the hot electron energy and temperature scaling relations under various laser-plasma conditionsMichel et al. (2013); Froula et al. (2012). One primary source of these hot electrons is two-plasmon decay (TPD), a major LPI in both directSeka et al. (2009); Turnbull et al. (2020) and indirectRegan et al. (2010) ICF schemes. In TPD, one laser wave decays into two electron plasma waves (EPW), efficiently transferring energy to the EPWsW.L.Kruer (2003), which can accelerate background electrons. In addition to experimental studies, scaling laws have been examined through first-principle particle-in-cell (PIC) simulationsCao et al. (2022); Cao and Ren (2023), considering both TPD and stimulated Raman scattering (SRS)—another key LPI that generates hot electronsRosenberg et al. (2018); Dewald et al. (2016). In SRS, a laser wave decays into an EPW and an electromagnetic wave (EMW)W.L.Kruer (2003).
@@
Various approaches have been developed to mitigate hot electron generation, including advancements in target designFollett et al. (2016) and broadening of laser bandwidthsYao et al. (2024); Lei et al. (2024); Wang et al. (2024); Liu et al. (2024); Ma et al. (2021); Follett et al. (2021); Gao et al. (2020); Zhao et al. (2017). Nevertheless, uncertainties persist regarding the angular distribution of hot electrons. Given hot electron energy and temperature, the impact of hot electrons depends on their angular characteristics. In conventional center hot-spot ignition for both directChristopherson et al. (2021) and indirect driveDewald et al. (2016) ICF, directional hot electrons that propagate towards the capsule center pose a greater preheating threat than those moving laterally. Similarly, the SI scheme benefits from directional hot electrons to augment the shock pressure. Conversely, in the emerging double-cone ignition (DCI) schemeZhang et al. (2020a), hot electrons generated in the compression cone and moving laterally may preheat the ignition cone, forming local pre-plasmas that reduce the energy coupling efficiency of the picosecond ignition laser pulse to the compressed fuelMacPhee et al. (2010); Li et al. (2013).
@@
Experiments have demonstrated that the angular distribution of hot electrons varies with laser illumination geometry and plasma conditions. Early studiesEbrahim et al. (1980) involving single-beam $CO_{2}$ laser pulses interacting with long-scale-length plasmas have revealed that hot electrons are concentrated at $45^{\circ}$ relative to the laser incident direction in the polarization plane. In contrast, more recent experimentsYaakobi et al. (2013) on the OMEGA laser facility have observed a wide divergence of hot electrons during implosions of spherical targets illuminated by 60 laser beams. In subsequent SI-relevant experimentsZhang et al. (2020b) on the OMEGA-EP laser facility, where intense single-beam laser pulses (with intensities of $5-10\times 10^{15}W/cm^{2}$) interacted with planar targets, the measured hot electrons associated with both TPD and SRSLi et al. (2020) exhibited directional characteristics. A more comprehensive experimental studyLiu et al. (2017) using single laser beam at varying intensities has shown that the angular distributions of hot electrons varies depending on the specific laser and target configurations. The broad angular spread of hot electrons is generally attributed to the wide EPW spectrum in the nonlinear phase of TPDMyatt et al. (2014). Nevertheless, the mechanisms behind the generation of hot electrons with various angular features remain insufficiently understood. A key factor influencing these angular characteristics is the laser beam geometry, as the daughter EPWs in TPD and SRS are primarily driven by wave vector matching across multiple laser beams in current ICF experimentsMyatt et al. (2014); Michel et al. (2015).
@@
In this article, we investigate the angular variations of TPD-generated hot electrons in relation to laser illumination geometry using fully kinetic PIC simulations and theoretical analysis. Our results show that laser geometry is the primary determinant of the angular features of hot electrons. In our study, we specifically focus on dual laser beams symmetric to the density gradient, while typical cases with asymmetric incidence are preliminarily analyzed using the theoretical model developed for symmetric configurations, with adjustments to account for asymmetry. There are two key reasons for emphasizing the symmetric dual beam cases. First, experiments have demonstrated that the overlapped intensity of laser beams in the same incident cone plays a dominant role in hot electron generationMichel et al. (2013); Regan et al. (2010). Our study using dual laser beams with identical incident angles is extendable to more beams within the same cone due to angular symmetry. Second, even though the total overlapped intensity in a cone matters most in hot electron generation, typical OMEGA-EP experiments have shown the dominance of dual-beam coupling rather than more beams in TPD growthMichel et al. (2012). We study incident angle $\theta_{in}$ ranging from $24^{\circ}$ to $55^{\circ}$, as $\theta_{in}\sim 60^{\circ}$ brings the laser reflection density near the quarter-critical density $n_{c}/4$, lowering the TPD threshold--- -a scenario already addressed in previous workLian et al. (2022); Zhou et al. (2023). Higher $\theta_{in}$ values shift the reflection density away from $n_{c}/4$ to lower density region, making TPD instability unlikely.
@@
This article is organized as follows. Section II describes the simulation setup. In Section III, we provide simulation results with theoretical explanations and discuss the underlying physics. Section IV offers further discussion on the applicability of our results to realistic experimental conditions. Section V presents the conclusions of this study.
@@ II Simulation setup
At time $t=0$, the p-polarized laser pulses are launched at the left boundary ($x=0$) of the simulation domain. For the simulations with $L_{n}=150\mu m$, the domain size is $L_{x}\times L_{y}=819.2c/\omega_{0}\times 3276.8c/\omega_{0}$ (or $43.5\mu m\times 173.8\mu m$) consisting of $4096\times 16384$ square cells with size $dx=dy=0.2c/\omega_{0}$ ($0.011\mu m$). Each cell is initially distributed 200 numerical particles (100 electrons, 50 fully ionized Carbon ions and 50 Hydrogen ions). It is worth noting that the cell size remains the same for all the simulations with varying domain sizes. The simulations progress with a time step of $dt=0.1414\omega_{0}^{-1}=0.025fs$ until $\sim 6ps$ when TPD has mostly reached a quasi-steady state. We employ open field boundaries and thermal particle boundaries along both x- and y-directions. During the simulations, we record all the hot electrons crossing the boundaries with energies exceeding 50 keV. The electromagnetic fields are diagnosed using the time Fourier transform (TFFT) moduleWen et al. (2019); Cao et al. (2022) with time windows of $\sim 1ps$. This module records the time Fourier-transformed fields throughout the entire space but only within specific frequency ranges: (0.45,0.55)$\omega_{0}$ for Langmuir waves, (0.95,1.05)$\omega_{0}$ for light waves. With frequency filtering, this method provides cleaner signals than the conventional spatial Fourier transform of the “snapshots" and can distinguish the directions of different wave modes. Of course, we still employ conventional “snapshots" diagnostics of all the particles and fields.
@@ III.1 Two major characteristics of TPD growth for varying $\theta_{in}$
The growth of TPD modes at different times can be demonstrated by $<\Delta n_{e}>(x,t)$ in figure (2). For $\theta_{in}=24^{\circ}$, TPD modes first grow near the $n_{c}/4$ density, then diffuse to $\sim 0.23n_{c}$ and reach the quasi-steady state after $\sim 3ps$. By contrast, TPD modes for $\theta_{in}=52^{\circ}$ initially grow in a much broader space range from $0.22n_{c}$ to $n_{c}/4$, and this range remains until TPD reaches a quasi-steady state.
@@ III.2 The excitation mechanism of side TPD modes for relatively larger $\theta_{in}$
For the two characteristics described in the previous section, the one corresponds to $\theta_{in}=24^{\circ}$ is driven through a forward-going common wave and has been well understoodZhang et al. (2014). Thus, we focus on the underlying physics of the side modes excitation for $\theta_{in}=52^{\circ}$. We find that the side EPWs are still common waves, but they are shared by TPD driven by one beam and SRS by the other, instead of two TPDs, so we name it TPD-SRS instability (TSI) in this article. This TSI mechanism can explain most major modes in figure (3b)(3d), especially those side modes with $k_{x}\sim 0$.
To better illustrate the TSI mechanism, we first recall the basic TPD and SRS theories of wave coupling and maximum growth conditions. As typical three-wave coupling processes, both TPD and SRS require matching conditions of frequency $\omega_{0,1,2}$ and wave vector $\vec{k}_{0,1,2}$ to be satisfied for the instability to grow:
@@
To provide clarity, we point out that the equation (7) of Rosenbluth gain was developed with WKB approximationsRosenbluth, White, and Liu (1973), which basically requires that all the involved EPWs and EMWs are not close to their critical densities. This condition, however, does not necessarily hold for the dominant wave modes in our theoretical [the peak in figure (5a)] and numerical [figure (4b)] results, where the scattered light of SRS is sufficiently close to its own critical density to violate the WKB approximationMichel (2023). This makes the SRS absolutely unstable and the calculated convective gain from the equation (7) no longer reflects the original meaning of the amplification factor of a delta-function seed. However, we still follow the Rosenbluth gain formula [equation (7)] here. First, except for those dominant modes, this gain formula works well for most matching modes at other densities and can properly reflect their gains. Second, the “erroneous" results of this formula for the dominant absolute SRS modes are actually overwritten by the effects of finite width of the laser beams, as we will explain in the following few paragraphs [equation (9) and figure (5b)].
Using Rosenbluth gain theory, the total gain of the TSI common waves driven simultaneously by TPD and SRS can be evaluated by considering the contributions from both TPD and SRS:
@@
In addition to the example discussed above, our theoretical predictions of the peak common wave locations align well with the PIC simulation results under various laser plasma conditions. Figure (6) presents the $k_{y}-n_{e}$ phase spaces of $E_{y}$ overlaid with the peak locations of $G_{tot}$ for the simulation (i) and (v) [table 1] with $\theta_{in}=52^{\circ}$. In the two $k_{y}-n_{e}$ figures, the modes near (a) $0.22n_{c}$ and (b) $0.23n_{c}$ with $k_{y}\sim 0.9\omega_{0}/c$ are the TSI common EPWs, while those at the same density but with much larger $k_{y}$ correspond to the other paired TPD modes. These figures show that the $G_{tot}$ peak locations consistently match the theoretical TSI common wave locations.
@@ III.3 The angular characteristics of hot electrons for different $\theta_{in}$
By comparing the common wave fields to the locations of the accelerated electrons shown in figure (8a) and (8b), it becomes evident that the hot electrons are concentrated in bands positioned between the peaks of the common wave fields. The parallel yellow bands along the y-direction suggest that these electrons are being accelerated by the wave fields and are moving along the x-direction. These findings indicate that the symmetric, divergent common waves can produce collimated hot electrons.
Figure 8: (a) The $E_{cx}$ fields [equation (III.3)] of two common waves and electron energy density $\epsilon_{e}$ [equation (14)] in the overlap region near $n_{e}=0.217n_{c}$ at $t=1.9ps$ for the PIC simulation (i) [Table 1] with $\theta_{in}=52^{\circ}$. The data in (a) are obtained by filtering the space spectra of the common waves out of the total electric field using Fourier and inverse-Fourier transforms.

### v2

@@ Abstract
In inertial confinement fusion (ICF) implosions, the preheating risks associated with hot electrons generated by laser plasma instabilities (LPI) are contingent upon the angular characteristics of these hot electrons for a given total energy. Using particle-in-cell simulations, we reveal a novel multi-beam collaborative mechanism of two-plasmon decay (TPD) and stimulated Raman scattering (SRS), and investigate the angular variations of hot electrons generated from this shared TPD-SRS (STS) instability driven collectively by dual laser beams with varying incident angles $\theta_{in}$ ($24^{\circ}$ to $55^{\circ}$ at the incident plane) for typical ICF conditions. In the simulations with $\theta_{in}\gtrsim 44^{\circ}$, STS emerges as the dominant mechanism responsible for hot electron generation, leading to a wide angular distribution of hot electrons that exhibit both pronounced divergent and collimated components. The common Langmuir wave associated with STS plays a crucial role in accelerating both components. By properly modeling the STS common wave gains, we establish scaling relations between these gains and the energies of collimated and divergent hot electrons. These relations reveal that the divergent hot electrons are more sensitive to variations in gain compared to the collimated electrons. Additionally, the calculated gains qualitatively predict the asymmetry in hot electron angular distributions when the density gradients deviate from the bisector of the laser beams. Our findings offers insights for hot electron generation with multiple beams, potentially complementing previous experiments that underscore the critical role of overlapped intensity from symmetric beams within the same cone and the dominance of dual-beam coupling.
@@ I Introduction
In inertial confinement fusion (ICF), a spherical capsule is imploded by multiple high-power laser beams to compress the central fuels to extreme density and pressure, thereby initiating self-sustained fusion reactionsAtzeni and Meyer-ter Vehn (2004). During the implosion, hot electrons (with energy above $\sim 50$ keV) produced by laser plasma instabilities(LPI)W.L.Kruer (2003) can preheat the uncompressed fuel and degrade the implosionChristopherson et al. (2021); Solodov et al. (2022). In most ICF schemesNuckolls et al. (1972); He et al. (2016); Zhang et al. (2020a); Lan (2022), such hot electrons need to be reduced to improve the implosion, except for shock ignitionBetti et al. (2007) (SI), in which hot electrons generated during the spike pulse can enhance the shock pressureBaton et al. (2020); Tentori et al. (2021) and enhance the compression. While ignition has been achieved using the indirect-drive scheme on the NIF laser facilityAbu-shawareb (2024) with hot electron threats minimized by unique experimental configurations, the pursuit of higher energy gainsBatani et al. (2023); Sui and Lan (2024) and alternative ignition schemesGopalaswamy et al. (2024); Williams et al. (2024); He et al. (2016); Zhang et al. (2020a); Lan (2022); Betti et al. (2007) still demands effective control and thorough characterization of hot electron effects. Thus, reliably characterizing the properties of hot electrons is crucial to evaluating their potential detrimental or beneficial impacts.
[…]
@@
Various approaches have been developed to mitigate hot electron generation, including advancements in target designFollett et al. (2016) and broadening of laser bandwidthsYao et al. (2024); Lei et al. (2024); Wang et al. (2024); Liu et al. (2024); Ma et al. (2021); Follett et al. (2021); Gao et al. (2020); Zhao et al. (2017). These studies primarily focus on the suppression of the total hot electron energy. Nevertheless, uncertainties remain concerning the angular distribution of hot electrons. Given hot electron energy and temperature, the impact of hot electrons depends on their angular characteristics. In conventional center hot-spot ignition for both directChristopherson et al. (2021) and indirect driveDewald et al. (2016) ICF, directional hot electrons that propagate towards the capsule center pose a greater preheating threat than those moving laterally. Similarly, the SI scheme benefits from directional hot electrons to augment the shock pressure. Conversely, in the emerging double-cone ignition (DCI) schemeZhang et al. (2020a), hot electrons generated in the compression cone and moving laterally may preheat the ignition cone, forming local pre-plasmas that reduce the energy coupling efficiency of the picosecond ignition laser pulse to the compressed fuelMacPhee et al. (2010); Li et al. (2013).
@@
Experiments have demonstrated that the angular distribution of hot electrons varies with laser illumination geometry and plasma conditions. Early studiesEbrahim et al. (1980) involving single-beam $CO_{2}$ laser pulses interacting with long-scale-length plasmas have revealed that hot electrons are concentrated at $45^{\circ}$ relative to the laser incident direction in the polarization plane. In contrast, more recent experimentsYaakobi et al. (2013) on the OMEGA laser facility have observed a wide divergence of hot electrons during implosions of spherical targets illuminated by 60 laser beams, which agrees well with the PIC simulationsYan et al. (2012, 2009). In subsequent SI-relevant experimentsZhang et al. (2020b) on the OMEGA-EP laser facility, where intense single-beam laser pulses (with intensities of $5-10\times 10^{15}W/cm^{2}$) interacted with planar targets, the measured hot electrons associated with both TPD and SRSLi et al. (2020) exhibited directional characteristics. A more comprehensive experimental studyLiu et al. (2017) using single laser beam at varying intensities has shown that the angular distributions of hot electrons varies depending on the specific laser and target configurations. The broad angular spread of hot electrons is generally attributed to the wide EPW spectrum in the nonlinear phase of TPDMyatt et al. (2014); Yan, Li, and Ren (2014), while the SRS hot electrons are considered more directionalYan, Li, and Ren (2014). Nevertheless, the mechanisms behind the generation of hot electrons with various angular features remain insufficiently understood. A key factor influencing these angular characteristics is the laser beam geometry, as the daughter EPWs in TPD and SRS are primarily driven by wave vector matching across multiple laser beams in current ICF experimentsMyatt et al. (2014); Michel et al. (2015).
@@
In this article, we investigate the angular variations of hot electrons in relation to laser illumination geometry using fully kinetic PIC simulations and theoretical analysis. Our results demonstrate that laser geometry is the primary determinant of the angular features of hot electrons, which are mainly driven by a novel shared TPD-SRS (STS) mechanism for moderate or high incident angles. These hot electrons exhibit both pronounced divergent and collimated components, differing from previous understanding, where SRS mainly produces collimated hot electrons and TPD generates widely spread hot electrons. In our study, we specifically focus on dual laser beams symmetric to the density gradient, as the theoretical model we developed under symmetric configurations is also capable of accounting for asymmetry. There are two key reasons for emphasizing the symmetric dual beam cases. First, these configurations can be extended to more beams within the same incident cone, where the overlapped intensity plays a dominant role in hot electron generationMichel et al. (2013); Regan et al. (2010). Second, typical OMEGA-EP experiments have shown the dominance of dual-beam coupling rather than more beams in TPD growthMichel et al. (2012). We study incident angle $\theta_{in}$ ranging from $24^{\circ}$ to $55^{\circ}$, as $\theta_{in}\sim 60^{\circ}$ brings the laser reflection density near the quarter-critical density $n_{c}/4$, lowering the TPD threshold--- -a scenario already addressed in previous workLian et al. (2022); Zhou et al. (2023). Higher $\theta_{in}$ values shift the reflection density away from $n_{c}/4$ to lower density region, making TPD instability unlikely.
@@
This article is organized as follows. Section II describes the simulation setup. In Section III, we provide simulation results with theoretical explanations and discuss the underlying physics. Section IV offers further discussion on the applicability of our results to realistic experimental conditions, Section V presents the conclusions of this study.
@@ II Simulation setup
At time $t=0$, the p-polarized laser pulses are launched at the left boundary ($x=0$) of the simulation domain. For the simulations with $L_{n}=150\mu m$, the domain size is $L_{x}\times L_{y}=819.2c/\omega_{0}\times 3276.8c/\omega_{0}$ (or $43.5\mu m\times 173.8\mu m$) consisting of $4096\times 16384$ square cells with size $dx=dy=0.2c/\omega_{0}$ ($0.011\mu m$). Each cell is initially distributed 200 numerical particles (100 electrons, 50 fully ionized Carbon ions and 50 Hydrogen ions). It is worth noting that the cell size remains the same for all the simulations with varying domain sizes. The simulations progress with a time step of $dt=0.1414\omega_{0}^{-1}=0.025fs$ until $\sim 6ps$ when TPD has mostly reached a quasi-steady state. We employ open field boundaries and thermal particle boundaries along both x- and y-directions. During the simulations, we record all the hot electrons crossing the boundaries with energies exceeding 50 keV. The electromagnetic fields are diagnosed using the time Fourier transform (TFFT) moduleWen et al. (2019); Cao et al. (2022) with time windows of $\sim 1ps$. This module records the time Fourier-transformed fields throughout the entire space but only within specific frequency ranges: (0.45,0.55)$\omega_{0}$ for Langmuir waves, (0.95,1.05)$\omega_{0}$ for light waves. With frequency filtering, this method provides cleaner signals than the conventional spatial Fourier transform of the “snapshots", which is also employed in our analysis, and can distinguish the directions of different wave modes.
@@ III.1 Two major characteristics of TPD growth for varying $\theta_{in}$
The growth of TPD modes at different times can be demonstrated by $<\Delta n_{e}>(x,t)$ in figure (2). For $\theta_{in}=24^{\circ}$, TPD modes first grow near $n_{c}/4$, then diffuse to $\sim 0.23n_{c}$ and reach the quasi-steady state after $\sim 3ps$. By contrast, TPD modes for $\theta_{in}=52^{\circ}$ initially grow in a much broader space range from $0.22n_{c}$ to $n_{c}/4$, and this range remains until TPD reaches a quasi-steady state.
@@ III.2 The excitation mechanism of side TPD modes for relatively larger $\theta_{in}$
For the two characteristics described in the previous section, the one corresponds to $\theta_{in}=24^{\circ}$ is driven through a forward-going common wave and has been well understoodZhang et al. (2014). Thus, we focus on the underlying physics of the side modes excitation for $\theta_{in}=52^{\circ}$. We find that the side EPWs are still common waves, but they are shared by TPD driven by one beam and SRS by the other, instead of two TPDs, so we name it shared TPD-SRS (STS) instability in this article. This STS mechanism can explain most major modes in figure (3b)(3d), especially those side modes with $k_{x}\sim 0$.
To better illustrate the STS mechanism, we first recall the basic TPD and SRS theories of wave coupling and maximum growth conditions. As typical three-wave coupling processes, both TPD and SRS require matching conditions of frequency $\omega_{0,1,2}$ and wave vector $\vec{k}_{0,1,2}$ to be satisfied for the instability to grow:
@@
Using Rosenbluth gain theory, the total gain of the STS common waves driven simultaneously by TPD and SRS can be evaluated by considering the contributions from both TPD and SRS:
@@
In addition to the example discussed above, our theoretical predictions of the peak common wave locations align well with the PIC simulation results under various laser plasma conditions. Figure (6) presents the $k_{y}-n_{e}$ phase spaces of $E_{y}$ overlaid with the peak locations of $G_{tot}$ for the simulation (i) and (v) [table 1] with $\theta_{in}=52^{\circ}$. In the two $k_{y}-n_{e}$ figures, the modes near (a) $0.22n_{c}$ and (b) $0.23n_{c}$ with $k_{y}\sim 0.9\omega_{0}/c$ are the STS common EPWs, while those at the same density but with much larger $k_{y}$ correspond to the other paired TPD modes. These figures show that the $G_{tot}$ peak locations consistently match the theoretical STS common wave locations.
@@ III.3 The angular characteristics of hot electrons for different $\theta_{in}$
By comparing the common wave field structure [figure (8a)] to the locations of the accelerated electrons [figure (8b)], using the dashed grids as a reference, it becomes evident that the hot electrons are concentrated in bands positioned between the peaks of the common wave fields. The parallel yellow bands along the y-direction suggest that these electrons are being accelerated by the wave fields and are moving along the x-direction. These findings indicate that the symmetric, divergent common waves can produce collimated hot electrons.
Figure 8: (a) The $E_{cx}$ fields [equation (III.3)] of two common waves and (b) electron energy density $\epsilon_{e}$ [equation (14)] in the overlap region near $n_{e}=0.217n_{c}$ at $t=1.9ps$ for the PIC simulation (i) [Table 1] with $\theta_{in}=52^{\circ}$. The dashed grids in the figures illustrate the alignment between the (a) $E_{cx}$ structures and (b) the accelerated electrons. The data in (a) are obtained by filtering the space spectra of the common waves out of the total electric field using Fourier and inverse-Fourier transforms.